### v1

# Statistics within UV–Visible Absorption Spectrum of Ethanolic Azobenzene

Abstract
We report a statistical simulation of the UV–visible absorption spectrum of trans-azobenzene in ethanol solution. Due to intermolecular interactions, the used explicit solvent environment necessitates accounting for numerous transitions for a spectrum covering the two energetically lowest lines, S1 and S2. Furthermore, the spectrum manifests vast variation as a function of the underlying local structure, in conjunction with previous observations for spectra of liquids in the X-ray regime. We disentangle the complex structure–spectrum relationship using a machine learning-based method known as the emulator-based component analysis. This structural decomposition outperforms commonly used principal component analysis in explained spectral variation and reveals a small subset of latent structural variables responsible for the total spectral variance. Among other structural characteristics, blueshifting of the S2 line occurs with fewer hydrogen bonds with the ethanol solvent, and a contracted N=N bond within the C–N=N–C bridge. The observed structural dependence of the absorption spectrum thus implies an overrepresentation of certain structural classes after a photoexcitation, potentially significant for the subsequent nuclear dynamics, photophysics, and photochemistry.
A liquid system exhibits a wealth of local atomistic structures, which is reflected as strong variation in their X-ray [1, 2, 3, 4, 5] and UV–visible [6, 7, 8] spectra. In general, sampling of the ensemble is required, as the effect is not necessarily captured by just a few representative structures and their respective spectra [9]. On one hand, statistical variation of the spectra implies pre-selection of structures for, e.g., absorption at a given wavelength. On the other hand, this dependence allows for interpreting the details of spectra for the purposes of characterization if the spectrally relevant structural degrees of freedom can be identified. In a liquid system, however, this is not completely straightforward. Fortunately, extensive simulations enable an in-depth study of the phenomenon when combined with analysis techniques capable of transforming numerical results into human-interpretable forms [10, 11, 12, 13].
We investigate statistical variation in the UV–visible spectrum of trans-azobenzene in ethanol solution using machine learning-based analysis of data based on sampling by molecular dynamics (MD). Photoisomerization between the trans and the cis forms [14] of the molecule is an archetypal topic of photochemistry and a mechanism utilized in applications [15, 16]. We sampled 30 000 structures from extended tight binding MD (xTBMD) using the GFN1-xTB method [17, 18] implemented in the CP2K software [19] v. 2024.2. In the simulations, a trans-azobenzene molecule was dissolved in 159 ethanol molecules in a cubic box of $L$ = 25.04 Å for the density 0.7840 g/cm3 interpolated for pure ethanol liquid at 300 K [20]. After equilibration of 20 ps, we sampled the canonical ensemble at 300 K with a 10 fs interval from a production run of 300 ps. We used a global Nosé–Hoover thermostat [21, 22] with chain length 3, a time constant $\tau$ of 1000 fs, and an MD timestep of 0.5 fs. For evaluation of spectra consisting of 100 transitions on each of the 30 000 structures, we applied time-dependent density functional theory (TD-DFT) implemented in ORCA v.6.0.0 [23]. We explicitly included ethanol within the cutoff radius of 3.0 Å, and simulated this cluster in a conductor-like polarizable continuum model [24]. We used the PBE exchange-correlation potential [25], the def2-SVP basis set [26] and def2/J auxiliary basis set [27] in the spectral simulations. We evaluated the cross section $\sigma_{i}$ of each transition $i$ for a randomly oriented sample using the transition matrix element in the velocity form, and then convolved this stick spectrum to differential cross section $\mathrm{d}\sigma/\mathrm{d}E$ in terms of energy $E$ by a Gaussian of 0.2 eV full width at half maximum. Finally, we transformed the differential cross section into wavelength $\lambda$ scale for $\mathrm{d}\sigma/{\mathrm{d}\lambda}$. We use a constant width in energy, because a constant width in wavelength makes the lines artificially broad for short wavelengths.
Figure 1: Results of the spectrum calculations and variance captured by different dimensionality reduction techniques. a: Simulated ensemble average UV–visible absorption spectrum for trans-azobenzene in ethanol, together with an experiment reproduced from Ref. 28. The blue-shaded area depicts the standard deviation of the data set, indicating significant statistical variation. The gray-shaded areas depict the two regions of interest (ROIs I and II) defined for the subsequent analysis. Computational results have been shifted $-50.6$ nm to match the experiment. Additionally, the figure presents an example local structure prepared with Jmol [29]. b: Spectral R2 score as a function of the rank of emulator-based component analysis (ECA) and principal component analysis (PCA), evaluated using the neural network emulator after the respective dimensionality reduction. The first two ECA components explain a large portion of the spectral ROI variance, while the higher components show rapidly diminishing R2 score gains. Structural decomposition by PCA covers significantly less spectral variance for the same rank. c: Structural R2 score as a function of decomposition rank. The ECA vectors cover a small fraction of the structural variance compared to structural PCA. The result indicates that explaining structural variation alone does not explain the spectral variation.
Figure 1a shows the simulated ensemble-average spectrum shifted by $-50.6$ nm in comparison to the experimental one by Nägele and coworkers [28]. We apply the molecular notation S1 and S2 for the characteristic features in this spectrum, but note that in our simulation for the gas phase in the optimized geometry, the S2 line already consists of three transitions. In liquid, we find the photoexcited states of trans-azobenzene to be mixed with states of the ethanol, which results in a wealth of possible final states to be excited – and to be simulated with explicit solvation. In our data, the lines S1 and S2 typically consist of several notably intense transitions and numerous weaker ones. The S1 feature (390 nm–520 nm) is almost unobservable, and we focus on the S2 peak (270 nm–390 nm). We observe significant statistical variation between the spectra of individual local structures, as shown by the standard deviation $\sigma$ in Figure 1a. For the following analysis, we define two regions of interest (ROI), split at the maximum of the S2 peak. Then, we study the mean differential cross section within each ROI as a function of the structure. We first train a neural network to predict the two z-score standardized ROI values for a structure, and then utilize this emulator in emulator-based component analysis (ECA) [10] for dimensionality reduction to identify the decisive structural characteristics behind the observed spectral variation.
We pre-process the structural information using a local version of the many-body tensor representation (LMBTR) [30] to obtain a vector $\mathbf{d}$ for each of the data points. We then apply z-score standardization for $\mathbf{\tilde{d}}$, the vector input to a neural network. The LMBTR descriptor consists of Gaussian-smeared element-wise radial distance distributions for multiple centers in the system. As these centers, we use all the nitrogen and carbon sites in the azobenzene molecule, and eight additional virtual centers located perpendicular to a plane defined by nearby atoms in the azobenzene molecule. The first four virtual centers lie ±$\pm 2$ from each nitrogen atom. The latter four centers lie ±$\pm 2$ perpendicular from the arithmetic mean position of the para and ortho carbon atoms of each ring. We ran a joint search [12] of the LMBTR and neural network hyperparameters, choosing the best-performing combination (for details, see SI). The final model had an R2 score (coefficient of determination) of 0.891 on a test set of 5 000 points, split from the data pool of 30 000 points prior to model selection, and thus previously unseen by the model. This performance is similar to that observed for X-ray spectra [5, 12, 13]. In addition to good ML and ECA performance, demonstrated for X-ray spectra, the LMBTR descriptor is also interpretable by a human.
The ECA algorithm aims to identify a few spectrally most relevant structural degrees of freedom, ordered by their spectral significance. We project the z-score standardized structural feature vectors $\mathbf{\tilde{d}}$ of the test set onto the basis vectors $\mathbf{v}_{i}$. Thus, each basis vector results in a single latent coordinate $t_{i}$ for a data point. Furthermore, an approximation
$\mathbf{\tilde{d}}^{(k)}=\sum_{i=1}^{k}t_{i}\mathbf{v}_{i}$ (1)
for each data point $\mathbf{\tilde{d}}$ is obtained in the given rank $k$. In the ECA procedure, vectors $\mathbf{v}_{i}$ are tuned rank-by-rank to maximize the explained spectral variance for the neural network emulator prediction over all data $\mathbf{\tilde{d}}^{(k)}$. We first optimize the basis vectors for dimensionality reduction using the training set, and then apply this dimensionality reduction in the analysis performed using the test set. The results of the analysis in the z-score standardized space can be converted to the original descriptor space by the corresponding inverse transform.
As seen in Figure 1b, rapid convergence towards the emulator performance limit is observed when emulation is carried out on data points projected onto the low-dimensional subspace. The ECA method is analogous to principal component analysis (PCA) with the key difference being that the optimization target of the former is the explained ROI intensity variance. In contrast, applying PCA to the structural data points $\mathbf{\tilde{d}}$ aims at finding the decisive components of structural variance, not necessarily reflected by the spectral response as seen in Figure 1b. Vectors from structural PCA cover significantly less spectral variance after the corresponding evaluation. Perhaps even more notably, the dominant structural degree of freedom is spectrally irrelevant. The different optimization goal of the methods is also evident in the structural R2 score as a function of the decomposition rank, presented in Figure 1c, which shows PCA to surpass the ECA decomposition. Thus, the majority of the spectral ROI intensity variance is explained by a minority of structural variance. We note that another method, using partial least squares fitting, has previously been compared to ECA, with the latter excelling in this comparison [10].
Next, we turn to the analysis of the first two ECA component vectors, that indicate the structural changes responsible for most spectral ROI intensity variation. We first analyze the spectral effect, and then the associated structural changes. We define two spectral properties: (i) ROI value difference (ROI I – ROI II), and (ii) total ROI value sum. These properties can be directly linked in the two latent coordinates $t_{1}$ and $t_{2}$, as indicated in Figure 2. The spectrally most significant structural degree of freedom $t_{1}$ corresponds to the ROI value difference (Figure 2a), whereas the second most significant degree of freedom $t_{2}$ captures the ROI value sum (Figure 2d). The two properties are practically independent of each other, illustrated by the low cross-correlation in Figures 2b and 2c. From Figures 2a and 2d we conclude that ROI I value is high when $t_{1}$ is high and $t_{2}$ is low. Similarly, ROI II is high when both $t_{1}$ and $t_{2}$ are low. Thus, the structural changes causing a blueshift of the S2 line are indicated by the latent coordinate $t_{1}$, as the effects from the second ECA component cancel out. For a structural interpretation of such effect, we calculate the difference 10 $\mathbf{v}_{1}$ for a well-expected change in this coordinate, and then transform the respective ECA component vector to the original descriptor space. A further division by the squared distance to a corresponding center yields site-wise and element-wise curves with radial behavior proportional to that of the radial distribution function.
Figure 2: The relationship between spectral properties and the latent coordinates from ECA using the test set with the corresponding Pearson’s correlation coefficients $r$. a: The difference between the two ROI values is highly correlated with the spectrally most significant structural variable $t_{1}$. b: The correlation between ROI value difference and the second most significant structural variable $t_{2}$ is weak. c: The total ROI value sum is weakly correlated with $t_{1}$, but d: highly correlated with $t_{2}$. For details, see text.
Figure 3 shows a selection of curves from the aforementioned comparison. The shorter-wavelength ROI I favors structures with less hydrogen bonding between the nitrogen atoms of trans-azobenzene and the OH group of the ethanol solvent. This is evident from the reduction of hydrogen and oxygen in the curves with an azobenzene N at the origin (Figure 3a,b). This is accompanied by an increase of carbon near the nitrogen center, which we interpret as rotation of the ethanol molecule. Moreover, a blueshift is indicative of shorter N–N and CN=Northo-carbon bond lengths. The simultaneous reduction of N–C curves in short ranges (Figures 3a,b) indicates a shift towards longer N–CN=N bond lengths at the azobenzene C–N=N–C bridge. We note that other notable structural changes are implied as well, and refer to the Supplementary Information for the complete LMBTR vectors. A minor effect may arise from, for example, bending of the N=N–C angles typically associated with the ”inversion pathway” of the photodynamics [15]. To confirm this, we calculated the Pearson correlation coefficient between the ROI values and the angles, finding a slight correlation of around $\sim$0.1–0.2. On the other hand, we note that the C–N=N–C dihedral, associated with the ”rotation pathway” [15], has negligible correlation with the ROI values.
We performed a similar analysis on the ROI value sum in terms of the corresponding latent coordinate $t_{2}$. The results, shown in the Supplementary Information, indicate that hydrogen bonding and the bond lengths have a much lesser effect. Instead, the ROI sum is affected by, for example, twisting of the molecule and an increase in the amount of the ethanol solvent near the very tips of the carbon rings. The result can be affected by a molecule being excluded from the explicitly treated cluster due to the necessitated cutoff distance, but is still indicative of a closer ethanol molecule.
Figure 3: Element-wise change in radial number density ($\Delta$RND) implied by a ten-unit rise of latent coordinate $t_{1}$ associated with ROI value difference from a: the nitrogen atoms, and from b: carbon atoms of the C-N=N-C bridge. We denote regions corresponding to the weakening of the hydrogen bonding and possible rotation of the ethanol molecules with an asterisk (*). Additionally, the triangles ($\triangledown$) denote regions indicating shortening of N=N or CN=N–C bonds, and the square ($\Box$) denotes a region indicating an increase in the length of the N–C bond. The curves shown are an average for the two N atoms or the two CN=N atoms.
We proceed with the discussion of the results. The observed strong statistical variation of the UV–visible spectrum highlights the importance of proper ensemble sampling for the analysis of spectra of liquid systems. In this case, the variation is similar, if not greater, to what has been reported for X-ray spectra of liquids [2, 4, 31, 12, 13]. Furthermore, our analysis shows that the spectrum is affected by a wide range of structural properties, as a multitude of features in the ECA curves were non-zero (see Figure 3 and Supplementary Information). In spite of the large number of involved features, we emphasize that this subspace still covers a relatively small fraction of the total structural variance (Figures 1b and 1c). Last, the fact that the ECA curves show a contribution from oxygen atoms is a direct indication that the ethanol solvent has a significant effect on the spectrum.
During preliminary studies, we found that relatively large neural networks without regularization outperformed smaller or regularized ones. Interestingly, funnel-shaped neural networks with a linearly decreasing number of neurons at each layer seemed to allow for sufficient complexity while using less resources than their rectangle-shaped counterparts of uniform width. These large networks had a tendency to heavily overfit to the training data while still generalizing well to the validation data and, eventually, the test data. This phenomenon is likely an example of a machine learning model reaching the interpolating regime with emergent self-regularization capability [32]. In contrast, the model selection in our previous works with X-ray spectra favored simpler architectures [5, 12, 13]. Furthermore, compared to the X-ray regime, a larger number of LMBTR centers with some virtual sites are necessary for sufficient machine learning performance for the UV–visible spectrum of ethanolic trans-azobenzene. While the structure–spectrum relation in the X-ray regime is not necessarily straightforward, these results are a strong indication of an even greater complexity of UV–visible spectra. The conclusion is reasonable in the sense that both orbitals involved in the transition matrix element extend over a large volume, whereas the core-orbital participating in an X-ray process is quite localized.
The presented analysis necessitated numerous approximations. For example, the cutoff value for the explicit solvent is an obvious one. We chose the cutoff to be as large as possible while still enabling the simulation of $\sim$104 structure–spectrum pairs. Additionally, we ignore the possible spectral effects arising from photodynamics and vibrational states. Despite this, the reasonability of our data is supported by the fact that the general shape of the spectrum agrees with the experiment. Regarding the analysis of this work, the challenges and limitations of the ECA method have been discussed in detail in Ref. 13. Most notably, the interpretation of the ECA results is dependent on the properties of the descriptor. Therefore, development of a new, highly interpretable, atomistic descriptor for machine learning can have a significant positive impact on the method in the future.
Instead of representing two specific atomistic structures, the curves in Figure 3 reflect shifts in the structural distributions between the two selected points in the latent coordinate space. While not necessarily always the case, for the UV–visible spectrum of ethanolic trans-azobenzene, the two latent coordinates link the respective structural classes to a shift of the S2 line, and reveal potentially important implications for photoprocesses in liquids. Namely, as a consequence of the observed spectral behaviour, the used excitation wavelength introduces a preference for certain initial structures for the following photodynamics, potentially observable by, e.g., pump–probe experiments. Solvent properties have also been found to have an effect on the photochemistry of azobenzene [15], which could partially be a manifestation of this pre-selection.
In conclusion, the UV–visible absorption spectrum of ethanolic trans-azobenzene shows a broad statistical variation analogous to that found in X-ray spectra. Our subsequent analysis condensed the spectral variance to a few structural degrees of freedom, enabling a partial interpretation of the complex structure–spectrum relationship. We found that photon excitation on the short-wavelength side of the S2 peak favors structures with, among other characteristics, weaker hydrogen bonding with the ethanol solvent, and contraction of the N=N bond. Thus, the chosen incident photon energy may favor certain structures for the following photodynamics.
Acknowledgements
We acknowledge CSC – IT Center for Science, Finland, for providing computing resources. We thank Dr. N. Runeberg at CSC for support with the Puhti supercomputer, and Prof. J. Wachtveitl for providing the experimental UV–visible absorption spectrum of trans-azobenzene in ethanol solution.
Supplementary Information
.1 Machine learning details
We split the data set of 30 000 structure–spectrum pairs into three parts: (i) a training set of 24 000 points for training the neural network and fitting the ECA vectors, (ii) a validation set of 1 000 points for early stopping of neural network training and (iii) a test set of 5 000 points for testing the model and ECA. For the performance metric we chose the coefficient of determination or R2 score which represents the spectral variance covered by the model. An R2 score of 1 refers to perfect prediction performance whereas a value of 0 refers to performance where the model always predicts the mean of the data set. We z-score standardized the LMBTR vectors and the ROI values before any machine learning tasks. This means that we subtracted the according training data mean from each of the vector components, after which we divided the values by their standard deviations in the training set.
As the machine learning model we chose a fully connected feed-forward neural network emulator implemented using PyTorch [^1] A. Paszke et. al. PyTorch: An Imperative Style, High-Performance Deep Learning Library. Advances in Neural Information Processing Systems 32, 8024–8035, 2019. url:https://dl.acm.org/doi/10.5555/3454287.3455008.. As illustrated in Figure 4, we used funnel-shaped neural networks with linearly decreasing number of neurons at each layer (excluding the input layer). Both the descriptor and the neural network have a wide range of tunable hyperparameters. Therefore, we ran a joint hyperparameter search [^2] E. A. Eronen, A. Vladyka, Ch. J. Sahle, and J. Niskanen. Structural descriptors and information extraction from X-ray emission spectra: aqueous sulfuric acid. Physical Chemistry Chemical Physics, 26:22752–22761, 2024. doi:10.1039/D4CP02454K. including both the descriptor and neural network hyperparameters as presented in Table 1. We trained the models to predict the z-score standardized ROI intensities in mini-batches of 512 points each. The training was stopped if the performance on the separate early stopping data set of 1 000 points did not improve in 10 consecutive epochs. We used the rectified linear unit activation function and the Adam optimizer with an initial learning rate of 10-3 to minimize the R2 loss (equals $1-\mathrm{R}^{2}$ score).
For selection of the best combination of hyperparameters, we used five-fold cross-validation using the aforementioned training method to evaluate the performance of each trial model. In the end, we chose the model with the best mean cross-validation R2 score and trained the final model using the full training data of 24000 data points.
For ECA we used the implementation by Vladyka et al. [^3] A. Vladyka, E. A. Eronen, and J. Niskanen, Implementation of the emulator-based component analysis, Journal of Computational Science 83, 102437 (2024). doi:10.1016/j.jocs.2024.102437. To ensure that a proper minimum is found, we fitted the each ECA vector 25 times with different initial guesses and chose the one with best performance before proceeding with the optimization of the next vector.
Figure 4: A schematic of a funnel-shaped neural network used in this work.
Table 1: The grid for the joint search of the LMBTR and neural network (NN) hyperparameters. The best performing combination is indicated in bold face.
Hyperparameter Grid points
LMBTR grid min, Å {0}
grid max, Å {4, 5, 6, 7}
number of grid points {10, 20, 30}
Gaussian width, Å {0.4, 0.6, 0.8, 1.0}
NN weight decay {0}
number of hidden layers {2, 3, 4, 5, 6}
width of the first hidden layer {1024, 2048, 4096, 8192, 16384}
.2 A schematic of all LMBTR centers
Figure 5 shows a schematic of all the centers that we included in the LMBTR descriptor. The notation is used throughout the analysis of the structure–spectrum relationship. In the main text, we show the average of element-wise curves centered at N0 and N1, and the average of element-wise curves centered at C5 and C8.
Figure 5: A schematic of the centers for the LMBTR descriptor of the azobenzene molecule. We chose all the nitrogen and carbon atoms of the molecule and a set of eight virtual points (V) as centers. The first four are located $\pm$2 Å off-plane from the nitrogen atoms, with the plane defined by three nearby atoms (N1, N0 and C5, or N0, N1 and C8). The latter four are located $\pm$2 Å off-plane from the carbon rings (R), with the plane defined by the two ortho-carbon atoms (C4 and C6, or C9 and C10) and a para-carbon atom (C2 or C13).
.3 ROI value difference with large overall intensity: full list of structural changes
Figure 6 shows the complete list of structural changes implied by a blueshift in the S2 peak, according to the applied analysis procedure. The component vector, originally in the z-score-standardized descriptor space, has been inverse transformed to the original descriptor space after which each feature has been divided by the squared distance to the corresponding center. As mentioned in the main text, several curves show behavior indicative of bending of the N=N–C angles. In addition to the results discussed in the main text, the curves at centers C2 and C13 also show a shortening of the C–C bonds. The virtual centers located off-plane either from the nitrogen atoms or the carbon rings show large values near zero. This is an artifact originating from the division by the squared distance to the center. We note that the curves also show other effects that are difficult to interpret; for example the hydrogen and carbon curves of the virtual centers.
Figure 6: The full LMBTR vector corresponding to the structural shift behind the increase of the ROI value difference. We denote regions corresponding to weakening of hydrogen bonds with an asterisk (*), shortening of N=N bond with triangles ($\triangledown$), shortening of or C–C bonds with diamonds ($\diamond$), increase in the N–C bond lengths with squares ($\Box$), and bending of the N=N–C angles with circles ($\circ$). Numerous other less interpretable factors also play a role in the structure–spectrum relationship.
.4 Structural changes implied by the increase of total ROI value
Similarly to the analysis of the ROI value difference, Figure 7 shows the complete list of structural changes implied by an increase in the total ROI value, i.e. the sum of the average differential cross section within each ROI. Again, several structural factors affect the spectrum, some being more interpretable than others. These include, for example, presence of ethanol solvent near the tips of the azobenzene rings. As is the case in the analysis of the ROI difference, twisting of the azobenzene molecule is also implied. In contrast to the ROI difference, changes in hydrogen bonding between the azobenzene molecule and the ethanol solvent do not have a notable effect on the total ROI value.
Figure 7: The full LMBTR vector corresponding to the structural shift behind the increase in total ROI value. We denote regions corresponding to bending of the N=N–C angles with circles ($\circ$) and increase in the amount of ethanol solvent near the tips of the azobenzene rings with plus signs (+). Many other less interpretable factors also play a role in the structure–spectrum relationship.